# Improving Rule-based Reasoning in LLMs via Neurosymbolic Representations

### v1

@@ Discussion
  • The Neurosymbolic LLM outperforms all other models on trained problems, while also not sacrificing performance on testing problems.
  • The Standard LLM performs well on simpler tasks but struggles with problems requiring intermediate reasoning or symbolic representation. The Standard LLM has a 85.59% higher loss and a 78.57 times lower score than the Neurosymbolic LLM
  • The CoT LLM’s reliance on multi-step reasoning introduces opportunities for errors, particularly in tasks involving non-trivial intermediate computations. The Standard LLM has an 89.08% higher loss and a 28.36 times lower score than the Neurosymbolic LLM
  • The LoRA LLM’s inability to generalize to unseen tasks underscores the advantage of neurosymbolic encoding for maintaining task flexibility. The Standard LLM has a 76.65% higher loss and a 20.65 times lower score than the Neurosymbolic LLM
These findings validate the utility of neurosymbolic encoding as a useful tool for enhancing the reasoning capabilities of LLMs, particularly in domains where precision and rule-following are required, while also providing insights into the model’s internal representations by converting hidden states into interpretable and compositional symbolic vectors.
@@ 6 Conclusion
We introduce a neurosymbolic method that bridges the strengths of LLMs and symbolic reasoning systems to address challenges in rule-based reasoning tasks. By encoding LLM hidden states into symbolic representations, solving problems in a symbolic domain, and merging solutions back into the LLM, our approach achieves significant improvements in mathematical reasoning tasks. Experimental results demonstrate superior accuracy and reliability compared to traditional methods like CoT reasoning and fine-tuning with LoRA modules.
@@ Acknowledgements
Appendix A Appendix
Appendix B Determining Problem Types and Intervention Thresholds
@@
For problems seen during training, we expect that $\mathbf{result}$ will be approximately equal to a problem type seen during training, since one of the encoders purposes is to represent the correct problem type in it’s neurosymbolic vector output. For problems not seen during training, the expected behavior is that $\mathbf{result}$ should be dissimilar to all problem types seen during training. This fact allows us to prevent the neurosymbolic system from intervening on untrained problems, which allows us to benefit from improved performance on trained problem types while not sacrificing performance on untrained problem types.
For example, if the LLM is asked ”What is $920$ mod $895$?”, the neurosymbolic vector generated by the encoder is queried for it’s problem type, and the dot product of this vector is taken with the neurosymbolic vector representing every problem type. For this problem, the various dot product similarities are shown in table 3. The table shows that the Modulo problem type has the highest similarity to the problem type queried from our neurosymbolic vector, which means that the system will use the program corresponding to modulo to generate the solution neurosymbolic vector.
@@
Figure 4 shows the distribution of dot product similarities of different problems, where the dot product similarity is the maximum dot product similarity of the queried problem type vector and all problem type vectors defined during training. This suggests that we can avoid intervention on problems not seen during training by imposing a maximum similarity threshold, by imposing that if the maximum dot product similarity between the queried problem type and problem types seen during training is below a threshold, then we do not use the neurosymbolic system and we do not intervene. In this situation, the output of the LLM would be it’s standard forward pass output. Figure 4 suggests this threshold should be $0.8$, which is what we use in this study.
@@
Appendix C Decoder Fine Tuning
@@
As mentioned in Section 4.4, the decoder network requires fine tuning to properly enhance the performance of the LLM on the rule-based reasoning tasks. This is because the decoder needs to learn how to insert information about the solution to the task into the LLMs forward pass in a way that is both effective and non-disruptive. Figures LABEL:loss_vs_epoch and LABEL:score_vs_epoch illustrate that as fine-tuning progresses, both cross-entropy loss decreases and task performance improves, highlighting the importance of optimizing the decoder within the LLM context to maximize performance.
Appendix D Error Analysis of Chain-of-Thought Reasoning
@@
D.1 Representation Errors
One source of errors is when the LLM incorrectly represents the input numbers, and consequently gets the wrong final answer. In the below example, we query the LLM with ”Solve the following problem step by step: What is $601$ plus $106$”. We can see that in the LLMs CoT response, it says $601$ has 6 hundreds and 1 tens and 1 ones, which is clearly incorrect, which leads to the LLM getting the wrong final answer.
@@ D.1 Representation Errors
D.2 Intermediate Step Errors
The CoT approach also underperforms the standard LLM in more complicated tasks, such as bitwise OR as well. One reason for this is because these problems require multiple rules or algorithms being applied, which increase the opportunity for the LLM to make an error. For example, if we query the LLM with ”Solve the following problem step by step: What is $513$ OR $107$”, the first step the LLM executes is to convert each number into binary. Afterwards, it will perform the bitwise OR on the bitstrings for each number, and convert the resultant bitstring back into a decimal number. The below example showcases a situation where the LLM incorrectly converts one of the input numbers from decimal into binary ($513$ is $1000000001$, not $1000001001$). This causes the final solution of the LLM using CoT to be incorrect.
@@ D.2 Intermediate Step Errors
D.3 Stuck in Infinite Loops
Another source of errors when using CoT is the LLM being stuck in an endless cycle. In the below example, we query the LLM with ”Solve the following problem step by step: What is $661$ plus $420$”. The LLM starts with the correct procedure, but keeps applying the procedure to digits that are not in the original problem. This process repeats endlessly, resulting in the LLM failing to respond with the correct answer.

### v2

@@ Discussion
  • The Neurosymbolic LLM outperforms all other models on trained problems, while also not significantly sacrificing performance on testing problems (i.e., Addition and Integer Division).
  • The Standard LLM performs well on simpler tasks but struggles with problems requiring intermediate reasoning or symbolic representation. The Standard LLM has a 87% higher loss and a 25.5 times lower score than the Neurosymbolic LLM.
  • The CoT LLM’s reliance on multi-step reasoning introduces opportunities for errors, particularly in tasks involving non-trivial intermediate computations. The CoT LLM has a 91% higher loss and a 16.9 times lower score than the Neurosymbolic LLM.
  • The LoRA LLM’s inability to generalize to unseen tasks underscores the advantage of neurosymbolic encoding for maintaining task flexibility. The LoRA LLM has a 86% higher loss and a 13.8 times lower score than the Neurosymbolic LLM.
These findings validate the utility of neurosymbolic encoding as a useful tool for enhancing the reasoning capabilities of LLMs, demonstrating an average of 88.6% lower cross entropy loss and 15.4 times more problems correctly solved than the baselines. The advantages of our method are evident particularly in domains where precision and rule-following are required, while also providing insights into the model’s internal representations by converting hidden states into interpretable and compositional symbolic vectors.
@@ 7 Conclusion
We introduce a neurosymbolic method that bridges the strengths of LLMs and symbolic reasoning systems to address challenges in rule-based reasoning tasks. By encoding LLM hidden states into neurosymbolic representations, solving problems in a symbolic domain, and merging solutions back into the LLM, our approach achieves significant improvements in mathematical reasoning tasks. Experimental results demonstrate superior accuracy and reliability compared to traditional methods like CoT reasoning and fine-tuning with LoRA modules.
@@ Acknowledgements
Appendix
Appendix A Determining Problem Types and Intervention Thresholds
@@
For problems seen during training, we expect that $\mathbf{result}$ will be approximately equal to a problem type seen during training, since one of the encoders purposes is to represent the correct problem type in its neurosymbolic vector output. For problems not seen during training, the expected behavior is that $\mathbf{result}$ should be dissimilar to all problem types seen during training. This fact allows us to prevent the neurosymbolic system from intervening on untrained problems, which allows us to benefit from improved performance on trained problem types while not sacrificing performance on untrained problem types.
For example, if the LLM is asked “What is $920$ mod $895$?”, the neurosymbolic vector generated by the encoder is queried for its problem type, and the dot product of this vector is taken with the neurosymbolic vector representing every problem type. For this problem, the various dot product similarities are shown in table 3. The table shows that the Modulo problem type has the highest similarity to the problem type queried from our neurosymbolic vector, which means that the system will use the program corresponding to modulo to generate the solution neurosymbolic vector.
@@
Figure 4 shows the distribution of dot product similarities of different problems, where the dot product similarity is the maximum dot product similarity of the queried problem type vector and all problem type vectors defined during training. This suggests that we can avoid intervention on problems not seen during training by imposing a maximum similarity threshold, by imposing that if the maximum dot product similarity between the queried problem type and problem types seen during training is below a threshold, then we do not use the neurosymbolic system and we do not intervene. In this situation, the output of the LLM would be its standard forward pass output. Figure 4 suggests this threshold should be $0.8$ because that marks the point of minimal overlap between the two distributions, which is what we use in this study.
@@
Appendix B Performance Comparison to Non-Mathematical Problems
As discussed in Section Discussion, a key limitation of the LoRA module (and, more broadly, standard supervised fine-tuning approaches) is its inability to generalize to unseen problem types (such as addition or integer division in this study). This arises because LoRA modules are always active during inference—they cannot be selectively deactivated when the model is queried with tasks unrelated to their training data. In contrast, the NS LLM dynamically determines whether to intervene, allowing it to skip symbolic execution for unfamiliar prompts. This selective gating prevents degradation in model performance on out-of-distribution inputs (Kalajdzievski, 2024).
To evaluate this property beyond held-out mathematical operations, we test the NS LLM on a set of non-mathematical questions. We design seven qualitatively different topic categories: (i) philosophy, (ii) ethics, (iii) history, (iv) psychology, (v) science fiction, (vi) technology, (vii) art and culture. Each category contains five diverse open-ended prompts. For instance, one question under philosophy is: ‘‘Does morality exist independently of humans?’’
To determine whether the NS LLM intervenes on such prompts, we compute the maximum dot product similarity between the encoder-generated neurosymbolic vector and each of the problem type vectors in the training vocabulary.
Figure 5: Histogram of maximum problem type similarity for training problems vs. non-mathematical queries. None of the non-math queries exceed the 0.8 threshold.
Figure 5 shows that the maximum similarity for all non-mathematical queries remains below the 0.8 threshold described in Section A. This confirms that the NS LLM correctly suppresses decoder intervention for out-of-distribution prompts.
These results demonstrate that the NS LLM preserves the LLM’s original behavior on non-mathematical tasks, avoiding performance degradation seen in fine-tuned systems like LoRA. The selective activation mechanism of our approach ensures compatibility with general-purpose LLM usage beyond symbolic arithmetic.
Appendix C Computational Complexity
We analyze the time and space requirements of three settings: (a) a vanilla Transformer, (b) Transformer inference with key–value caching as implemented in LLaMA models, and (c) our neurosymbolic extension that inserts an encoder–symbolic–decoder block at layer $\ell^{\star}$.
Notation.
$n$ — sequence length seen by a layer during training, $d$ — model hidden state width ($4096$), $L$ — number of layers ($32$), $v$ — VSA dimensionality ($2048$), $D$ — maximal digit length encoded ($5$), $p$ — number of different problem types ($10$)
C.1 Baseline Transformer
During training every layer the vanilla transformer computes full self–attention and a two‑layer feed‑forward network. Following the derivation in Vaswani et al. (2017), the asymptotic per forward pass cost is
$\text{Time}_{\text{train}}$ $=\mathcal{O}\!\bigl{(}L\,\bigl{(}n^{2}d+n\,d^{2}\bigr{)}\bigr{)},$ (10)
$\text{Space}_{\text{train}}$ $=\mathcal{O}(L\,n\,d)+\Theta(\#\text{LLM params}),$ (11)
where the first term stores activations for back‑propagation and the second term is the fixed parameter memory. Without caching, inference has the same complexity but omits the activations (Pope et al., 2023; Dao et al., 2022).
C.2 Transformer Inference with KV Caching
LLaMA style decoding stores past key–value pairs, so a fresh token attends to $n$ cached tokens but does not recompute the $n^{2}$ attention matrix:
$\text{Time}_{\text{KV}}$ $=\mathcal{O}\!\bigl{(}L\,(n\,d+d^{2})\bigr{)}\quad\text{per new token},$ (12)
$\text{Space}_{\text{KV}}$ $=\mathcal{O}(L\,n\,d)+\Theta(\#\text{LLM params}),$ (13)
yielding linear rather than quadratic scaling in $n$ at inference time (Grattafiori et al., 2024; Shazeer, 2019; Ainslie et al., 2023).
C.3 Neurosymbolic Extension
At layer $\ell^{\star}$ we add:
(i) a dense encoder $W_{\!e}\!\in\!\mathbb{R}^{d\times v}$, (ii) symbolic computation in a VSA of width $v$, (iii) a dense decoder $W_{\!d}\!\in\!\mathbb{R}^{v\times d}$.
Encoder/decoder cost.
Each is a single matrix–vector product per token, giving $\mathcal{O}(d\,v)$ real operations.
Neurosymbolic cost.
Binding and unbinding use FFT‑based circular convolution; an FFT of length $v$ costs $\Theta(v\log v)$ operations (Cooley & Tukey, 1965).
The neurosymbolic procedure starts by decoding each digit of each number by binding with the inverse vectors of each possible digit value (i.e., 1’s place, 10’s place, …, $10^{D}$’s place), then binding the results with the inverse vectors of each possible base 10 number (i.e., 0 through 9). Thus, to decode the two numbers, $10\times D$ ($\Theta(D)$) FFT operations are performed.
In addition to querying the digit value of the two numbers from the output of the encoder, the problem type is also queried (this determines which mathematical operation is executed on the two input numbers). Querying the decoders involves first binding with the inverse of the problem type tag, followed by binding with the each problem type (e.g., modulo, multiplication, etc.). This results in $p+1$ FFT operations being performed to determine the problem type.
Once the problem type and input numbers are determined, the mathematical operation corresponding to the highest similarity queried problem type is executed. In our study, we considered ten such operations. Below is the run time of each operation in python:
All runtimes below follow the implementation details of CPython’s PyLong big‑integer arithmetic (Python Software Foundation, 2024).
  • Addition / Subtraction
    Algorithm:
    schoolbook carry chain
    Runtime: $\Theta(D)$
  • Multiplication
    Algorithm chosen by CPython 3.11 :
    Operand size Method & complexity
    $<70$ limbs [^3] A limb is one machine word of the internal base $\beta=2^{30}$ used by CPython’s big‑integer (“PyLong”) representation. Roughly, $70$ limbs correspond to $70\times 30\approx 2100$ bits, i.e. $\approx 630$ decimal digits. grade‑school (schoolbook) $\mathcal{O}(D^{2})$
    70707070240240240240 limbs Karatsuba $\mathcal{O}(D^{1.585})$
[…]
    $\geq 2^{14}$ limbs Schönhage–Strassen FFT $\mathcal{O}\!\bigl{(}D\log D\log\log D\bigr{)}$
    Overall worst‑case bound used in our analysis:
    $M(D)=\Theta\!\bigl{(}D\log D\log\log D\bigr{)}.$
  • Integer Division
    Algorithm:
    Burnikel–Ziegler + Newton reciprocal
    Runtime: $\Theta\!\bigl{(}M(D)\bigr{)}$
  • Modulo
    Algorithm:
    same core routine as division
    Runtime: $\Theta\!\bigl{(}M(D)\bigr{)}$
  • GCD
    Algorithm:
    Lehmer + binary Euclidean
    Runtime: $\Theta\!\bigl{(}M(D)\,\log D\bigr{)}$
  • Square mod ($x^{2}\bmod y$)
    Algorithm: one multiplication + one modulo
    Runtime: $\Theta\!\bigl{(}M(D)\bigr{)}$
  • LCM ($\operatorname{lcm}(x,y)\bmod 1000$)
    Algorithm: $\,\tfrac{xy}{\gcd(x,y)}$
    Runtime: $\Theta\!\bigl{(}M(D)\,\log D\bigr{)}$
  • Bitwise AND / OR / XOR
    Algorithm:
    limb‑wise bit operations
    Runtime: $\Theta(D)$
Total symbolic overhead.
For one token the neurosymbolic block executes
$\underbrace{\mathcal{O}(d\,v)}_{\text{encoder+decoder}}\;+\;\underbrace{\mathcal{O}\!\bigl{(}(10D+p+1)\,v\log v\bigr{)}}_{\text{digit \& problem–type queries}}\;+\;\underbrace{\mathcal{O}\!\bigl{(}M(D)\,\log D\bigr{)}}_{\text{worst‑case arithmetic}},$
so that
${\;\text{Time}_{NS}=\Theta\!\bigl{(}d\,v\;+\;(D+p)\,v\log v\;+\;M(D)\log D\bigr{)}\;}\quad\text{per forward pass at layer }\ell^{\star}.$
In our experiments $D=5$ and $p=10$ are fixed constants, and $M(D)\log D$ is likewise constant, giving the practical bound
$\Theta\!\bigl{(}d\,v+v\log v\bigr{)}\;\;=\;\;\Theta\!\bigl{(}v\,(d+\log v)\bigr{)}.$
Because this cost is independent of $n$ and $L$, while the key–value–cached Transformer in  (12) scales as $\mathcal{O}\!\bigl{(}L(nd+d^{2})\bigr{)}$, the standard KV path asymptotically dominates whenever $v<d$ (which holds for our setting, $v=2048<d=4096$). Thus the neurosymbolic block adds only a small, constant overhead to each decoder step.
Space complexity.
The space overhead of the neurosymbolic procedure at inference time consists of the encoder and decoder parameter matrices ($2dv$) and a temporary symbolic vector of size $v$. Thus, the total is
$\text{Space}_{NS}=\Theta(dv),$
which is asymptotically smaller than the key–value cache baseline in (13), which scales as $\mathcal{O}(Lnd)$. Since $v\ll nL$ in practical settings, and the number of LLM parameters far exceeds that of the encoder and decoder, the symbolic block contributes only a negligible additional memory footprint.
Appendix D Mixing Ratio Ablations
In this study, we use a 50/50 weighted sum to combine the neurosymbolic decoder output with the LLM hidden state at the intervention layer. While this approach is simple and effective, alternative mechanisms exist for fusing two streams of information.
One such alternative is the use of RMS Layer Normalization, which is commonly employed in the residual pathways of large transformer models, including LLaMA 3.1 (Grattafiori et al., 2024). In this scheme, the sum of the hidden state and decoder output is normalized by the root mean square of its elements and scaled by trainable parameters.
To evaluate whether RMS normalization improves performance over the 50/50 mixing strategy, we run an ablation study comparing the two approaches across several arithmetic tasks. Results are shown in Table 4. Overall, the 50/50 mixing method achieves superior accuracy and lower loss on nearly every task. These findings support our use of a fixed mixing ratio as the preferred integration mechanism in the NS LLM architecture.
Table 4: Performance of NS LLM using 50/50 mixing vs. RMS Layer Normalization.
Problem Type 50/50 Score 50/50 Loss RMS Score RMS Loss
Addition 98.7 0.093 98.6 0.140
Division 97.4 0.066 96.1 0.210
Multiplication 95.6 0.314 95.1 0.399
Modulo 98.7 0.093 97.4 0.277
GCD 94.2 0.205 88.4 0.459
LCM 87.3 1.051 81.0 1.441
Square Mod 58.9 2.818 56.1 3.189
Bitwise AND 91.2 0.755 92.3 0.809
Bitwise XOR 99.4 0.094 97.8 0.270
Bitwise OR 97.6 0.093 88.4 0.422
Appendix E Decoder Fine Tuning
@@
As mentioned in Section 4.4, the decoder network requires fine tuning to properly enhance the performance of the LLM on the rule-based reasoning tasks. This is because the decoder needs to learn how to insert information about the solution to the task into the LLMs forward pass in a way that is both effective and non-disruptive. Figures LABEL:loss_vs_step and LABEL:score_vs_step illustrate that as fine-tuning progresses, both cross-entropy loss decreases and task performance improves, highlighting the importance of optimizing the decoder within the LLM context to maximize performance. Here, one fine-tuning step corresponds to processing a single batch, with one backpropagation performed per batch.
Appendix F Error Analysis of Chain-of-Thought Reasoning
@@
F.1 Representation Errors
One source of errors is when the LLM incorrectly represents the input numbers, and consequently gets the wrong final answer. In the below example, we query the LLM with “Solve the following problem step by step: What is $601$ plus $106$”. We can see that in the LLMs CoT response, it says $601$ has 6 hundreds and 1 tens and 1 ones, which is clearly incorrect, which leads to the LLM getting the wrong final answer.
@@ F.1 Representation Errors
F.2 Intermediate Step Errors
The CoT approach also underperforms the standard LLM in more complicated tasks, such as bitwise OR as well. One reason for this is because these problems require multiple rules or algorithms being applied, which increase the opportunity for the LLM to make an error. For example, if we query the LLM with “Solve the following problem step by step: What is $513$ OR $107$”, the first step the LLM executes is to convert each number into binary. Afterwards, it will perform the bitwise OR on the bitstrings for each number, and convert the resultant bitstring back into a decimal number. The below example showcases a situation where the LLM incorrectly converts one of the input numbers from decimal into binary ($513$ is $1000000001$, not $1000001001$). This causes the final solution of the LLM using CoT to be incorrect.
@@ F.2 Intermediate Step Errors
F.3 Stuck in Infinite Loops
Another source of errors when using CoT is the LLM being stuck in an endless cycle. In the below example, we query the LLM with “Solve the following problem step by step: What is $661$ plus $420$”. The LLM starts with the correct procedure, but keeps applying the procedure to digits that are not in the original problem. This process repeats endlessly, resulting in the LLM failing to respond with the correct answer.